suppose the length of a rectangle is 2x + 1 and the width is 3x-2 . find the value of x if the perimeter is equal to 38 inches

Answer by ikleyn(52800)

.
suppose the length of a rectangle is 2x + 1 and the width is 3x-2 . find the value of x if the perimeter is equal to 38 inches
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

2L + 2W = 38, or

2*(2x+1) + 2*(3x-2) = 38,  --->

4x + 2 + 6x - 4 = 38,

10x - 2 = 38  --->  10x = 38 + 2  --->  10x = 40  --->  x = 40/10 = 4.

Answer.  x = 4.
         The length = 2x+1 = 9.  The width = 3x-2 = 10.
         The width is longer than the length.
         Author !?